OpenStudy (anonymous):

Anybody able to solve 25=-4.9t^2 + 23.4t +1.3? pleasee?

OpenStudy (anonymous):

ick!

OpenStudy (anonymous):

its too hard to factor and the quadratic forula doesnt seem to work..

OpenStudy (anonymous):

The quadratic formula should work. What happens when you try it?

OpenStudy (anonymous):

maybe start with \[0 = -4.9 t^2+23.4 t-23.7\] and then \[49t^2-234t+237=0\] and then the quadratic formula

OpenStudy (anonymous):

quadratic formula is a formula . it always works

OpenStudy (anonymous):

what would you then do with the roots you get ? i got 0.86 and -5.6 mabye my math is wrong?

OpenStudy (anonymous):

Those can be valid roots. Just because they're not integers doesn't mean they're not valid roots.

OpenStudy (anonymous):

or if you just need a numerical answer cheat even easier http://www.wolframalpha.com/input/?i=25%3D-4.9t^2+%2B+23.4t+%2B1.3

OpenStudy (anonymous):

Also, if this is for a story problem where t represents time, you can usually throw out the negative root and just use the positive root. Depends on the problem.

OpenStudy (anonymous):

the original question was how long is the baseball above 25 meters, with the equation h=-4.9t^2 +23.4t +1.3 if that makes any more sense . h=height in m, t =time

OpenStudy (anonymous):

yeah that is fine. between the two zeros it is above 25

OpenStudy (anonymous):

so the zeros i got make sense?

OpenStudy (anonymous):

from about 1.46 second to about 3.32 seconds or if it really asks "how long" then 3.32-1.46 whatever that is

OpenStudy (anonymous):

@ladybug you have the right idea but your zeros are wrong

OpenStudy (anonymous):

alright so i just have to redo my math?

OpenStudy (anonymous):

The zeros you got could make sense, but don't appear to with the story given. Try breaking it down again with the quadratic formula. For 25=-4.9t^2 + 23.4t +1.3, You'll want to egt rid of the 25 first, so it should be 0=-4.9t^2 + 23.4t -23.7. A = -4.9 B = 23.4 C = -23.7

OpenStudy (anonymous):

yes i did do that! somwhere from then after i messed up?

OpenStudy (anonymous):

Show us your next step and we'll go through it here just to be sure.

OpenStudy (anonymous):

-23=+/- sqr root of 23.4^2 -4(-4.9)(23.7)

OpenStudy (anonymous):

just a suggestion that it makes life easier to deal with whole numbers and positive leading coefficient (not that it really matters) and i would solve \[49t^2-234t+237=0\]

OpenStudy (anonymous):

then add the decimals in after? because this is sosmthing im handing in for marks lol

OpenStudy (anonymous):

no no, don't put them back in. just do the calculation with whole numbers might be easier is all

OpenStudy (anonymous):

oh okay!

OpenStudy (anonymous):

because you have something set equal zero, ten times it is zero as well. you will get the same answer as you would working with decimals, it just might be easier to use whole numbers. your choice

OpenStudy (anonymous):

oh i see makes sence thankyou!